Two 2.5 kg bowling balls are 0.50 m apart. What is the force of gravity of the first bowling ball due to the second? Compare this to the force of gravity on the second bowling ball due to the first.
1
Expert's answer
2020-12-10T11:02:34-0500

According to the third Newton's law, "When one body exerts a force on a second body, the second body simultaneously exerts a force equal in magnitude and opposite in direction on the first body."

Therefore, we may calculate the modulus of the gravitational force according to Newton's law of gravitation.


"F_{1,2} = G\cdot\dfrac{m_1m_2}{r^2} = 6.67\cdot10^{-11}\,\mathrm{N\cdot m^2/kg^2}\cdot\dfrac{2.5\,\mathrm{kg}\cdot2.5\,\mathrm{kg}}{(0.5\,\mathrm{m})^2} = 1.67\cdot10^{-9}\,\mathrm{N}."